OpenStudy (anonymous):

The number of bacteria in a certain culture doubles every hour. If there are 30 bacteria present in the culture originally i) How many bacteria will be present at the end of 2nd hour. ii) How many bacteria will be present at the end of nth hour.

OpenStudy (aravindg):

its a tricky qn

OpenStudy (anonymous):

1 hour 30 x 2 = 60 2 hour 60 x 2 = 120 And so on

OpenStudy (aravindg):

u jest need to apply ur brains

OpenStudy (anonymous):

B(t) = 30*2^t, where is the number of hours. B(1) = 30*2^1 = 60 after the first hour. B(2) = 30*2^2 = 120 after the second hour. B(n) = 30*2^n after the nth hour.